# Parameter-Efficient Subspace Optimization for LLM Fine-Tuning

### v1

@@ 2.2 Connection to Existing Benchmarks
Projected subspace.
A simple way to define a memory-efficient subspace is through low-rank projection, where $\mathcal{M}_{k}:R\in\mathbb{R}^{r\times n}\mapsto P_{k}R$ is parameterized by a left-projection matrix $P_{k}\in\mathbb{R}^{m\times r}$. This formulation can be extended to right-sided or two-sided projections. By applying the chain rule to $\nabla_{R}\ell(\widetilde{W}_{k}+P_{k}R)$, one obtains the projected subspace schemes analyzed in [he2024subspace, kozak2019stochastic, liang2024memory, zhao2024galore]; see Appendix C for details. Within PESO, GaLore [zhao2024galore], APOLLO [zhu2024apollo], and stochastic subspace descent [kozak2019stochastic] correspond to a restart strategy by reassigning $P_{k}$, while online subspace descent [liang2024memory] adopts a warm-start update of $P_{k}$ via online PCA.
Low-rank subspace.
The LoRA family defines the subspace $\{A_{k}B+AB_{k}:A\!\in\!\mathbb{R}^{m\times r},\,B\!\in\!\mathbb{R}^{r\times n}\}$, where the adapters $(A,B)$ jointly serve as both $\xi$ and the parameterization of $\mathcal{M}_{k}$. Consequently, a single Adam update of $(A,B)$ simultaneously updates the subspace and its coordinates, effectively realizing a warm-start scheme with $K=1$. Many LoRA variants can be viewed as modifications of this generic template: LoRA-Pro [wang2024lora] applies a different preconditioner, and PiSSA [meng2024pissa], LoRA-GA [wang2024loraga], and LoRA-One [zhang2025one] adjust initialization strategies, while other works modify learning rates or scaling rules. Our framework unifies these designs by interpreting them as specific choices of Opt or initialization within the same subspace structure.
SVD subspace.
A principled way to extract low-dimensional structure from matrices (such as $\Delta W$) is through Singular Value Decomposition (SVD), leading to the representation $\mathcal{M}_{k}:\lambda\in\mathbb{R}^{r}\mapsto U\mathrm{diag}(\lambda)V$. Here, $(U,V)$ define the subspace (exploration), while $\lambda$ is the low-dimensional coordinates (exploitation). This separation fits directly into Algorithm 1, enabling flexible optimization strategies for $(U,V)$ and $\lambda$. AdaLoRA [zhang2023adalora] exemplifies this parameterization, and our framework clarifies the roles of $(P_{k},\Lambda_{k},Q_{k})$ in their notation. We build on this principle in Section 3, where we propose a practical SVD-based variant PESO-LoRA-T.
Together, these three subspace categories illustrate how PESO unifies existing PEFT methods under a single framework, setting the stage for designing new practical algorithms and proving convergence in the following sections.
@@ 3 PESO-LoRA: Practical Algorithms from the Framework
The previous section establishes a conceptual link between PEFT methods and classical subspace minimization, providing a unifying interpretation. Building on this perspective, we now develop a concrete algorithm PESO-LoRA, which extends LoRA using guidance from our PESO framework. We present two variants: PESO-LoRA-R leverages a full gradient restart strategy to improve exploration of subspaces, and PESO-LoRA-T is a SVD-based method that enhances exploitation through more effective optimization within each subspace.
@@ 3.1 Full Gradient Restart
We now introduce an important variant of the UpdateSubspace subroutine in the restart category (see (6)) that enables convergence to stationarity in the full-parameter space. We design UpdateSubspace so that each new subspace $\mathcal{S}_{k}$ induced by $\mathcal{M}_{k}$ remains well aligned with the current full gradient $G_{k}$. We call this scheme full gradient restart:
@@
Thus, each restart ensures that $\mathcal{S}_{k}$ captures information from full gradients, with initial progress comparable to a standard SGD step. In the literature on subspace methods, incorporating the full gradient into $\{\mathcal{S}_{k}\}$ is critical for convergence guarantees [conn1994iterated, zhang2025scalable]. In particular, one can show that $\|\nabla_{W}\ell\|\to 0$ provided that $G_{k}:=\nabla_{W}\ell(W_{k})$ lies in $\mathcal{S}_{k}$. Building on this, we demonstrate in Section 5 that full gradient restart ensures convergence to a stationary point of the original problem (1) by interleaving projected steepest-descent steps with subspace updates.
@@
7:   $(U_{k},\Lambda_{k},V_{k})\leftarrow{\tt SVD}(-G_{k})$. $\triangleright$ Top-$r$ SVD of $G_{k}$
8:   Set $A_{k-1}\leftarrow\frac{1}{\sqrt{\gamma}}U_{k}\sqrt{\Lambda_{k}}$ and $B_{k-1}\leftarrow\frac{1}{\sqrt{\gamma}}\sqrt{\Lambda_{k}}V_{k}$.
@@
10:  $(A_{k},B_{k})\leftarrow\texttt{AdamW}(A_{k-1},B_{k-1})$. $\triangleright$ One AdamW step on $(A_{k-1},B_{k-1})$
@@
A practical construction of $\mathcal{S}_{k}^{\text{FG}}$ is to compute a rank-$r$ SVD of $G_{k}$ and define the subspace as the span of its top singular directions. This ensures that $\mathcal{S}_{k}^{\text{FG}}$ captures the main structure of $G_{k}$, while the approximation error $\|G_{k}-P_{\mathcal{S}_{k}^{\text{FG}}}(G_{k})\|$ is governed by the spectral tail of $G_{k}$. Crucially, this tail is independent of the rank gap in the objective, underscoring a key distinction between representation deficiency (e.g., LoRA) and update efficiency. In practice, given $(U_{k},V_{k})$ from the rank-$r$ SVD of $G_{k}$, one can also restart with $\mathcal{M}_{k}(\xi_{k}^{\text{new}})=-\eta_{k}U_{k}V_{k}$, which corresponds to the update from the recent training benchmark Muon [jordan2024muon], providing improved stability over (8).
One practical advantage of full gradient restart is that it acts as a “plug-and-play” mechanism for existing PEFT methods. It can be applied with a moderate frequency $K$ to reduce the cost of SVD while still guiding subspace exploration effectively. Recent work, such as GaLore [zhao2024galore] and LoRA-One [zhang2025one], has demonstrated the empirical benefits of leveraging the full gradient. In particular, the recent variants LoRA-GA [wang2024loraga] and LoRA-One [zhang2025one] can be interpreted as special cases of LoRA with full gradient restart applied only at initialization. To achieve convergence in the full-parameter space, we propose PESO-LoRA-R (Algorithm 2), which embeds LoRA with a repeated restart mechanism every $K$ iterations. A detailed version of the pseudocode can be found in Appendix E.
To implement Algorithm 2, directly assigning $(A_{k-1},B_{k-1})\leftarrow 1/{\sqrt{\gamma}}(U_{k}\sqrt{\Lambda_{k}},\sqrt{\Lambda_{k}}V_{k})$ can cause instability due to mismatches in optimization states. A similar instability issue has also been reported for GaLore [chen2024fira]. For robustness, we propose alignment techniques to maintain consistency of subspace bases, momentum, and velocity. In particular, restarts often produce gradients with much larger magnitudes, leaving the Adam velocity “too cold” and causing unstable steps. We therefore apply a lightweight velocity alignment, rescaling the velocity to match the new gradient magnitude, $v\leftarrow\|g\|^{2}/\|v\|\,v$, together with a short $\beta_{2}$ warm-up, which we find to be the most important component for stable performance in practice; see more details in Appendix D. Finally, we remark both empirical evidence and theoretical results suggest that gradients $G_{k}$ in deep learning often have strong low-rank structure, making them especially suitable for efficient SVD-based approximations [cosson2023low, yang2023spectral, zhao2024galore].
@@ 3.2 Exploitation via SVD Subspace
6:   $(U_{k},V_{k})\leftarrow\texttt{AdamW}(U_{k-1},V_{k-1})$. $\triangleright$ One AdamW step on $(U_{k-1},V_{k-1})$
@@
8:  $\xi_{k}\leftarrow\texttt{AdamW}(\xi_{k-1})$. $\triangleright$ One AdamW step on $\xi_{k-1}$
@@ 4 Experiments
In this section, we conduct experiments to evaluate our methods across diverse tasks and models, comparing with standard LoRA-based approaches and full fine-tuning. We first assess natural language understanding on the GLUE benchmark [wang2018glue] by fine-tuning T5-base [raffel2020exploring]. We then evaluate natural language generation on Llama-2-7B [touvron2023llama] for tasks including mathematical reasoning, code generation, and general instruction tuning. Finally, we demonstrate that LoRA-PESO-R remains effective even under strict memory constraints when trained for more epochs. Implementation details are provided in Appendix F.
4.1 Natural Language Understanding Tasks
We fine-tune the T5-base model on a subset of GLUE, including MNLI, SST-2, CoLA, QNLI, and MRPC, and evaluate performance using test accuracy (%). Following the setting in [zhang2025one], we compare our method against several LoRA variants, including LoRA [hu2022lora], LoRA+ [hayou2024lora+], P-LoRA [zhang2024riemannian], PiSSA [meng2024pissa], LoRA-GA [wang2024loraga], LoRA-Pro [wang2024lora], and LoRA-One [zhang2025one]. For fairness, hyperparameters are tuned individually for each method.
The results are summarized in Table 2. PESO-LoRA-R and PESO-LoRA-T achieve the best performance on three of the five GLUE tasks (MNLI, SST-2, and QNLI), which are also the larger datasets. On the remaining tasks, PESO-LoRA-T ranks second. This demonstrates the overall efficiency and robustness of our approaches, with advantages most evident on larger datasets that demand longer training and stronger exploration–exploitation. Moreover, PESO-LoRA-T generally outperforms PESO-LoRA-R, but at the cost of $1.4\times$ more computation time, whereas PESO-LoRA-R runs at nearly the same speed as standard LoRA. Memory costs are comparable across all methods, so the choice ultimately depends on whether performance or efficiency is prioritized.
Table 2: Performance of fine-tuned T5-base on natural language understanding tasks with rank set to 8. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies excluding PESO-LoRA-T, which is shaded in gray and omitted from direct comparison due to its longer runtime.
@@ 4.2 Natural Language Generation Tasks
Following prior work [wang2024loraga, zhang2025one], we fine-tune the Llama-2-7B model on three datasets and evaluate on the corresponding downstream tasks. For mathematical reasoning, we use a 100k subset of MetaMathQA [yu2023metamath] and evaluate on GSM8K [cobbe2021training]. For general instruction tuning, we fine-tune on Alpaca [alpaca] and evaluate on MMLU [hendrycks2020measuring]. For code generation, we use a 100k subset of Code-Feedback [zheng2024opencodeinterpreter] and evaluate on HumanEval [chen2021evaluating], reporting PASS@1. To ensure fairness, all datasets are preprocessed to exclude overlaps with test sets. Results are shown in Table 3. Remarkably, our methods outperform baselines on two of the three tasks—mathematical reasoning and code generation—both involving larger training datasets, highlighting the substantial gains enabled by subspace exploration and exploitation in handling complex tasks.
For the more advanced model, we run the same three tasks by fine-tuning the Llama-3.1-8B [dubey2024llama] model, comparing the LoRA-based methods with PESO-LoRA-R, and report the results in Table 4. We find that on a more capable model, our method consistently outperforms all LoRA variants across all benchmarks.
Additionally, we compare our method with several pretraining-oriented approaches on PEFT tasks. Following the experimental setup in [zhu2024apollo], we evaluate PESO-LoRA-R against GaLore [zhao2024galore], Fira [chen2024fira], and APOLLO [zhu2024apollo] by fine-tuning Llama-3-8B-Instruct [llama3modelcard] on the Alpaca-en-demo dataset from LlamaFactory [zheng2024llamafactory] with rank $r=8$, and report the results on MMLU subtasks. As shown in Table 5, our PEFT-specialized method PESO-LoRA-R consistently outperforms these pretraining-oriented baselines in the fine-tuning regime.
We further conduct ablation studies by varying the rank $r$ and the restart frequency $K$. The results, reported in Appendix F.3, show that PESO-LoRA-R does not require a large rank to achieve strong performance and exhibits robust behavior across a wide range of restart frequencies. Memory and runtime comparisons are also provided in Appendix F.3, showing that PESO-LoRA-R has similar memory footprint as vanilla LoRA, and that their overall computational costs are comparable.
[…]
Table 4: Performance of fine-tuned Llama-3.1-8B on natural language generation tasks with rank set to 8. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies.
@@
Table 5: Comparison with pretraining-oriented subspace methods. We fine-tune Llama-3-8B-Instruct on Alpaca-en-demo with rank $r=8$ and evaluate on MMLU subtasks. Results are reported as accuracy (%). For GaLore, Fira, and APOLLO, we use the results reported in [zhu2024apollo]. Following their convention, we report our method’s average accuracy, using the average over 3 runs. Bold and underline indicate the highest and second-highest accuracies.

### v2

@@ 2.2 Connection to Existing Benchmarks
Projected subspace. A simple way to define a memory-efficient subspace is through low-rank projection, where $\mathcal{M}_{k}:R\in\mathbb{R}^{r\times n}\mapsto P_{k}R$ is parameterized by a left-projection matrix $P_{k}\in\mathbb{R}^{m\times r}$. This formulation can be extended to right-sided or two-sided projections. By applying the chain rule to $\nabla_{R}\ell(\widetilde{W}_{k}+P_{k}R)$, one obtains the projected subspace schemes analyzed in (He et al., 2024; Kozak et al., 2019; Liang et al., 2024; Zhao et al., 2024); see Appendix C for details. Within PESO, GaLore (Zhao et al., 2024), APOLLO (Zhu et al., 2024), and stochastic subspace descent (Kozak et al., 2019) correspond to a restart strategy by reassigning $P_{k}$, while online subspace descent (Liang et al., 2024) adopts a warm-start update of $P_{k}$ via online PCA.
Low-rank subspace. The LoRA family defines the subspace $\{A_{k}B+AB_{k}:A\!\in\!\mathbb{R}^{m\times r},\,B\!\in\!\mathbb{R}^{r\times n}\}$, where the adapters $(A,B)$ jointly serve as both $\xi$ and the parameterization of $\mathcal{M}_{k}$. Consequently, a single Adam update of $(A,B)$ simultaneously updates the subspace and its coordinates, effectively realizing a warm-start scheme with $K=1$. Many LoRA variants can be viewed as modifications of this generic template: LoRA-Pro (Wang et al., 2024b) applies a different preconditioner, and PiSSA (Meng et al., 2024), LoRA-GA (Wang et al., 2024a), and LoRA-One (Zhang et al., 2025b) adjust initialization strategies, while other works modify learning rates or scaling rules. Our framework unifies these designs by interpreting them as specific choices of Opt or initialization within the same subspace structure.
SVD subspace. A principled way to extract low-dimensional structure from matrices (such as $\Delta W$) is through Singular Value Decomposition (SVD), leading to the representation $\mathcal{M}_{k}:\lambda\in\mathbb{R}^{r}\mapsto U\mathrm{diag}(\lambda)V$. Here, $(U,V)$ define the subspace (exploration), while $\lambda$ is the low-dimensional coordinates (exploitation). This separation fits directly into Algorithm 1, enabling flexible optimization strategies for $(U,V)$ and $\lambda$. AdaLoRA (Zhang et al., 2023) exemplifies this parameterization, and our framework clarifies the roles of $(P_{k},\Lambda_{k},Q_{k})$ in their notation. We build on this principle in Section 3, where we propose a practical SVD-based variant PESO-LoRA-T.
Together, these three subspace categories illustrate how PESO unifies existing PEFT methods under a single framework. This perspective guides the design of practical algorithms with new exploration and exploitation techniques for PEFT, and it sets up the convergence analysis in the sections that follow.
@@ 3 PESO-LoRA: Practical Algorithms from the Framework
The previous section establishes a conceptual link between PEFT methods and classical subspace minimization, providing a unifying interpretation. Building on this view, we now develop a concrete algorithm PESO-LoRA, which combines PESO’s optimization principles with LoRA’s empirically effective PEFT parameterization. We present two variants: PESO-LoRA-R leverages a full gradient restart strategy to improve exploration of subspaces, and PESO-LoRA-T is a SVD-based method that enhances exploitation through more effective optimization within each subspace.
@@ 3.1 Full Gradient Restart
We now introduce an important variant of the UpdateM subroutine in the restart category (see (6)) that enables convergence to stationarity in the full-parameter space. We design UpdateM so that each new subspace $\mathcal{S}_{k}$ induced by $\mathcal{M}_{k}$ remains well aligned with the current full gradient $G_{k}$. We call this scheme full gradient restart:
@@
Thus, each restart ensures that $\mathcal{S}_{k}$ captures information from full gradients, with initial progress comparable to a standard SGD step. In the literature on subspace methods, incorporating the full gradient into $\{\mathcal{S}_{k}\}$ is critical for convergence guarantees (Conn et al., 1994; Zhang, 2025). In particular, one can show that $\|\nabla_{W}\ell\|\to 0$ provided that $G_{k}:=\nabla_{W}\ell(W_{k})$ lies in $\mathcal{S}_{k}$. Building on this, we demonstrate in Section 5 that full gradient restart ensures convergence to a stationary point of the original problem (1) by interleaving projected steepest-descent steps with subspace updates.
@@
7:   $(U_{k},\Lambda_{k},V_{k})\leftarrow{\tt SVD}(-G_{k})$.
8:   $A_{k-1}\leftarrow\frac{1}{\sqrt{\gamma}}U_{k}\sqrt{\Lambda_{k}}$, $B_{k-1}\leftarrow\frac{1}{\sqrt{\gamma}}\sqrt{\Lambda_{k}}V_{k}$.
@@
10:  $(A_{k},B_{k})\leftarrow\texttt{AdamW}(A_{k-1},B_{k-1})$.
@@
A practical construction of $\mathcal{S}_{k}^{\text{FG}}$ is to compute a rank-$r$ SVD of $G_{k}$ and define the subspace as the span of its top singular directions. This ensures that $\mathcal{S}_{k}^{\text{FG}}$ captures the main structure of $G_{k}$, while the approximation error $\|G_{k}-P_{\mathcal{S}_{k}^{\text{FG}}}(G_{k})\|$ is governed by the spectral tail of $G_{k}$. Crucially, this tail is independent of the rank gap in the objective, underscoring a key distinction between representation deficiency (e.g., LoRA) and update efficiency. In practice, given $(U_{k},V_{k})$ from the rank-$r$ SVD of $G_{k}$, one can also restart with $\mathcal{M}_{k}(\xi_{k}^{\text{new}})=-\eta_{k}U_{k}V_{k}$, which corresponds to the update from the recent training benchmark Muon (Jordan et al., 2024), providing improved stability over (8).
[…]
To implement Algorithm 2, directly assigning $(A_{k-1},B_{k-1})\leftarrow 1/{\sqrt{\gamma}}(U_{k}\sqrt{\Lambda_{k}},\sqrt{\Lambda_{k}}V_{k})$ can cause instability due to mismatches in optimization states. A similar instability issue has also been reported for GaLore (Chen et al., 2024). For robustness, we propose alignment techniques to maintain consistency of subspace bases, momentum, and velocity. In particular, restarts often produce gradients with much larger magnitudes, leaving the Adam velocity “too cold” and causing unstable steps. We therefore apply a lightweight velocity alignment, rescaling the velocity to match the new gradient magnitude, $v\leftarrow\|g\|^{2}/\|v\|\,v$, together with a short $\beta_{2}$ warm-up, which we find to be the most important component for stable performance in practice; see more details in Appendix D. Finally, we remark both empirical evidence and theoretical results suggest that gradients $G_{k}$ in deep learning often have strong low-rank structure, making them especially suitable for efficient SVD-based approximations (Cosson et al., 2023; Yang et al., 2023; Zhao et al., 2024).
@@ 3.2 Exploitation via SVD Subspace
6:   $(U_{k},V_{k})\leftarrow\texttt{AdamW}(U_{k-1},V_{k-1})$.
@@
8:  $\xi_{k}\leftarrow\texttt{AdamW}(\xi_{k-1})$.
@@ 4 Experiments
In this section, we conduct experiments to evaluate our methods across diverse tasks and models, comparing with standard LoRA-based approaches and full fine-tuning. We first assess natural language understanding on the GLUE benchmark (Wang et al., 2018) by fine-tuning T5-base (Raffel et al., 2020). We then evaluate natural language generation on Llama models for tasks including mathematical reasoning, code generation, and general instruction tuning. Finally, we demonstrate that LoRA-PESO-R remains effective even under strict memory constraints when trained for more epochs. Implementation details are provided in Appendix F.
Table 2: Fine-tuned T5-base on natural language understanding tasks with rank set to 8. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies excluding PESO-LoRA-T, which is shaded in gray and omitted from direct comparison due to its longer runtime.
@@
4.1 Natural Language Understanding Tasks
We fine-tune the T5-base model on a subset of GLUE, including MNLI, SST-2, CoLA, QNLI, and MRPC, and evaluate performance using test accuracy (%). Following the setting in (Zhang et al., 2025b), we compare our method against several LoRA variants, including LoRA (Hu et al., 2022), LoRA+ (Hayou et al., 2024), P-LoRA (Zhang and Pilanci, 2024), PiSSA (Meng et al., 2024), LoRA-GA (Wang et al., 2024a), LoRA-Pro (Wang et al., 2024b), and LoRA-One (Zhang et al., 2025b). For fairness, hyperparameters are tuned individually for each method.
The results are summarized in Table 2. PESO-LoRA-R and PESO-LoRA-T achieve the best performance on three of the five tasks, which also have larger datasets. On the remaining tasks, PESO-LoRA-T ranks second. This demonstrates the overall efficiency and robustness of our approaches, with advantages most evident on larger datasets that demand longer training and stronger exploration–exploitation. Moreover, PESO-LoRA-T generally outperforms PESO-LoRA-R, but at the cost of $1.4\times$ more computation time, whereas PESO-LoRA-R runs at nearly the same speed as standard LoRA. Memory costs are comparable across all methods, so the choice ultimately depends on whether performance or efficiency is prioritized.
@@ 4.2 Natural Language Generation Tasks
We evaluate our method on the more capable Llama-3.1-8B (Dubey et al., 2024) model across three downstream tasks. For mathematical reasoning, we fine-tune on a 100k subset of MetaMathQA (Yu et al., 2023) and evaluate on GSM8K (Cobbe et al., 2021). For general instruction tuning, we fine-tune on Alpaca (Taori et al., 2023) and evaluate on MMLU (Hendrycks et al., 2020). For code generation, we use a 100k subset of Code-Feedback (Zheng et al., 2024) and evaluate on HumanEval (Chen et al., 2021), reporting PASS@1. To ensure fairness, all datasets are preprocessed to exclude overlaps with test sets.
The results are reported in Table 3. PESO-LoRA-R consistently outperforms all LoRA variants across all benchmarks on this advanced model. Notably, the gains are most pronounced in mathematical reasoning and code generation—tasks which involves larger fine-tuning datasets, highlighting the substantial advantages of subspace exploration and exploitation for complex tasks. For completeness, Table 7 in Appendix F.2 reports results on Llama-2-7B (Touvron et al., 2023), evaluating both PESO-LoRA-T and PESO-LoRA-R, where our methods similarly achieve competitive performance against established baselines. We further compare against pretraining-oriented approaches, including GaLore (Zhao et al., 2024), Fira (Chen et al., 2024), and APOLLO (Zhu et al., 2024); Table 5 shows that our PEFT-specialized PESO-LoRA-R consistently outperforms these pretraining baselines in the fine-tuning regime.
We conduct ablations over the rank $r$ and restart frequency $K$. Results in Appendix F.3 show that PESO-LoRA-R achieves strong performance with small ranks and remains robust across a wide range of $K$. Appendix F.3 also reports memory and runtime comparisons, showing a memory footprint similar to vanilla LoRA and comparable overall computational cost. In particular, Table 11 shows that the additional SVD cost over LoRA is only 1.9% of total runtime in our experiments, indicating negligible overhead.
Table 3: Fine-tuned Llama-3.1-8B on natural language generation tasks with rank set to 8. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies.
@@
Table 4: Fine-tuned T5-base (4 epochs) on the SST-2 dataset. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies.
Method Epoch 1 Epoch 2 Epoch 3 Epoch 4
Table 5: Comparison with pretraining-oriented subspace methods. Fine-tuned Llama-3-8B-Instruct (Alpaca-en-demo, $r=8$). Results are MMLU accuracy (%). Baseline results from Zhu et al. (2024). Bold and underline indicate the highest and second-highest accuracies.